OpenStudy (anonymous):

Brennan has been playing a game where he can create towns and help his empire expand. Each town he has allows him to create 1.15 times as many villagers. The game gave Brennan five villagers to start with. Help Brennan expand his empire by solving for how many villagers he can create with 15 towns. Then explain to Brennan how to create an equation to predict the number of villagers for any number of towns. Show your work and use complete sentences.

OpenStudy (anonymous):

If he has 5 villagers, then he could multiply 5 by 1.15 as many times as the number of towns. Because of the Commutative property of Multiplication you can multiply any of those numbers in any order. So for simplicity using an exponent would be the best way to model the function. An = 5(1.15)^(n-1) Using this formula he can calculate the amount of villagers for any amount of towns. N represents the number of towns, just substitute N and solve. For 15 towns A15 = 5(1.15)^(15-1) A15 = 5(1.15)^14 A15 = 5(7.08) A15 = 35.4 In his case the range is restricted to positive integers so you have to round A15 to 35. I literally did this question on tuesday on my algebra 2 test for FLVS.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

If you copy I don't suggest you do it word for word... Just incase lol